OpenStudy (anonymous):

What is the solution to the equation 8 - 2b = -2/3(12b + 6) ?

OpenStudy (anonymous):

start by multiplying each side by 3 to get rid of the denominator

OpenStudy (anonymous):

@Acoolhottbabe

OpenStudy (gorv):

step 1 multiply both side by 3

OpenStudy (anonymous):

both sides of the equal sign?

OpenStudy (anonymous):

@jrain35 @gorv

OpenStudy (anonymous):

yes, so you should get 24 -6b=2(12b+6)

OpenStudy (gorv):

\[3*(3-2b)=3*\frac{ -2 }{ 3 }(12b+6)\]

OpenStudy (gorv):

sorry that is 8-2b not 3-2b

OpenStudy (anonymous):

okay that's what I got

OpenStudy (anonymous):

and its okay haha @gorv

OpenStudy (gorv):

on right side 3 will cancel 3 agree???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So after that do I add the parentheses?

OpenStudy (anonymous):

you cant add the parentheses, you need to simplify all terms then add common terms

OpenStudy (gorv):

multiply the number inside that is outside the paranthesis

OpenStudy (anonymous):

so distribute the -2 through the parentheses

OpenStudy (gorv):

3*8-3*2b=-2*12b-2*6

OpenStudy (gorv):

yes do it and tell me what you u got ...?? btw you are goin right way

OpenStudy (anonymous):

24 - 6 = (-24 + -12) ?? I'm not sure if that's correct.